Find the finite area between the curves x=y^2 (for y≥0), y=−2x + 10 and the y-axis
Answer #1

283-21410 :) : ( 1) 2 Point of intersection 일 25로 10g yty -10 = 이 2845g-4100 20 (944552)=o (3-2) (29+5) 2 =들 갸2=4 뜻 a (-2 21